NvidiaIntely (nvidiaintely):

Red mercury (II) oxide decomposes to form mercury metal and oxygen gas according to the following equation: 2HgO (s) \(\rightarrow\) 2Hg (l) + O2 (g). If 3.55 moles of HgO decompose to form 1.54 moles of O2 and 618 g of Hg, what is the percent yield of this reaction?

OpenStudy (photon336):

first let's make sure that the equation is balanced \[2HgO_{s}~\rightarrow 2Hg~+~O_{2}\ make sure that the number of atoms of each element are the same on both the product and reactant side. do you believe this is balanced and if so why?

OpenStudy (photon336):

FYI this is a decomposition reaction we start with one reactant and get two new products.

OpenStudy (photon336):

@NvidiaIntely once you've told me this we will continue

NvidiaIntely (nvidiaintely):

Okay, well it's not balanced.

OpenStudy (photon336):

let's see so we have 2 atoms of Hg and 2 atoms of oxygen on the reactant side. agree?

NvidiaIntely (nvidiaintely):

No, becuase there on the reactant side it says: 2HgO, and on the product side it says: 2Hg + O2

NvidiaIntely (nvidiaintely):

Which means, that there is 1 oxygen atom on the reactant side, and 2 on the product side.

OpenStudy (photon336):

@NvidiaIntely do you see that 2 in front of HgO?

OpenStudy (photon336):

to find the total atoms you would multiply that by each atom in the compound so 2*HgO would be 2 atoms of Hg and 2 atoms of Oxygen

NvidiaIntely (nvidiaintely):

yeah, you mean the apply's to the gO too?

NvidiaIntely (nvidiaintely):

Okay got yah.

OpenStudy (photon336):

yeah

OpenStudy (photon336):

so our equation is balanced now we get the molar masses of all the compounds HgO = 216.59 g/mol Hg = 200.59 g/mol O2 = 32 g/mol

OpenStudy (photon336):

you understand how to find molar masses right?

NvidiaIntely (nvidiaintely):

Not fully, I would like you to explain further if you have time.

OpenStudy (photon336):

please watch this 3 minute video on how to find the molar mass of a compound https://www.youtube.com/watch?v=F9NkYSKJifs

NvidiaIntely (nvidiaintely):

Okay, thanks!

NvidiaIntely (nvidiaintely):

Now how do I find the percent yield of this reaction?

OpenStudy (photon336):

we need to use the molar ratios i.e. the numbers in front of the molecules in the balanced equation.

OpenStudy (photon336):

\[\frac{ Hg }{ HgO }*(3.55~mol~HgO)\]

OpenStudy (photon336):

we would multiply the number of moles of HgO by the molar ratio to find out how many moles of each substance are produced.

NvidiaIntely (nvidiaintely):

Hoe do I find the molar ratio?

OpenStudy (photon336):

from the balanced equation, see the numbers in front of the compounds?

NvidiaIntely (nvidiaintely):

Yes.

OpenStudy (photon336):

watch this 2 minute video on stoichiometry https://www.youtube.com/watch?v=eM_n1HzYlJo

NvidiaIntely (nvidiaintely):

Okay.

OpenStudy (photon336):

so we're given 3.55 moles of HgO

OpenStudy (photon336):

we need to find the number of moles of each product that are produced

OpenStudy (photon336):

so we need to start with 3.55 moles of HgO

NvidiaIntely (nvidiaintely):

Yeah

OpenStudy (photon336):

say if we wanted to find the number of moles of moles of oxygen that should be produced

OpenStudy (photon336):

we multiply the number of moles of HgO by the molar ratio between HgO and O2 ensuring that O2 is in the numerator. in our balanced equation, O2 we've got 1 mole of that and 2 moles of HgO So O_{2} goes in the numerator and HgO in the denominator \[3.55~moles~HgO*(\frac{ O_{2} }{ 2HgO }) = 1.78~mole~O_{2}\]

NvidiaIntely (nvidiaintely):

Okay

OpenStudy (photon336):

now that's based off of the stiochiometry

OpenStudy (photon336):

now we should get 1.78 moles of O2 but we only produced 1.54 moles

NvidiaIntely (nvidiaintely):

Okay

OpenStudy (photon336):

so to find the % yield it's \[\frac{ No~moles~you~got~from~reaction }{ No~moles~from~stiochiometry }*100 = percent~~yield\]

OpenStudy (photon336):

Here is the % yield of for oxygen \[\frac{ 1.54 }{ 1.78}*100 = 88~percent\] now try the same thing for Hg

NvidiaIntely (nvidiaintely):

Okay, let me think.

NvidiaIntely (nvidiaintely):

13.2%

OpenStudy (photon336):

Show your work

OpenStudy (photon336):

First we multiply the number of moles by the molar ratio of the two \[3.55~moles~HgO~*\frac{ Hg }{ HgO } = 3.55~moles~Hg\]

NvidiaIntely (nvidiaintely):

Okay

OpenStudy (photon336):

now we convert the grams of Hg produced to moles \[618~grams~Hg*(\frac{ 1~mole~Hg }{ 200.59~grams } )= 3.0~moles~Hg\]

OpenStudy (photon336):

putting it all together we get the following: \[\frac{ 3.00 }{ 3.55 }*(100) = 87~percent \]

NvidiaIntely (nvidiaintely):

oh, okay....:/

NvidiaIntely (nvidiaintely):

But on my answer sheet/page whatever, it only shows: 13.2% 42.5% 56.6% 86.5%

OpenStudy (photon336):

I got something like 86.7% i may have rounded up somewhere

NvidiaIntely (nvidiaintely):

So then I should use 86.5% right.

OpenStudy (photon336):

yes

NvidiaIntely (nvidiaintely):

Okay thanks!

NvidiaIntely (nvidiaintely):

Also thanks @Photon336 for making my problem sovling SS go to 25!

OpenStudy (photon336):

haha alright no problem